OpenStudy (anonymous):

Square root of x+3^10

OpenStudy (danjs):

\[\sqrt{x}+ 3^{10}\] is what you wrote but i will retriceume \[\sqrt{x+3^{10}}\] right

OpenStudy (anonymous):

Thanks. Yes

OpenStudy (freckles):

wait we can't say assume ?

OpenStudy (danjs):

There really is nothing you can do there, except expand out 3^10

OpenStudy (freckles):

oh we can you purposely wrote that

OpenStudy (danjs):

hahah i dont know if i mistyped that or what

OpenStudy (freckles):

are you sure the question is \[\sqrt{x+3^{10}}\]?

OpenStudy (danjs):

that is what i retriceume, i mean assumed above

OpenStudy (freckles):

\[\text{ maybe this: } \sqrt{(x+3)^{10}} ?\]

OpenStudy (anonymous):

Correct

OpenStudy (freckles):

well this definitely not the same as \[\sqrt{x+3^{10}}\] but okay \[\sqrt{(x+3)^{10}}=((x+3)^{10})^{\frac{1}{2}}\]

OpenStudy (freckles):

do you know what 10(1/2)=?

OpenStudy (alexandervonhumboldt2):

then (x+3)^5

OpenStudy (anonymous):

Right. \[\sqrt{x+3^10=x+3^5?}\]

OpenStudy (anonymous):

Wasn't trying to extend the symbol throughout the simplified answer